OpenStudy (anonymous):

Could someone help me learn how to do this?? I will give a medal and fan to whoever help me learn the best! What is the simplified form of the expression sqrt (16c^4) - sqrt (c^2) + 3sqrt (c^2) + sqrt(9c^2)

OpenStudy (anonymous):

@reemii @sfozturk61 @wmckinely

OpenStudy (zzr0ck3r):

sqrt(a*b) = sqrt(a)*sqrt(b) is the whole problem

OpenStudy (zzr0ck3r):

and sqrt(c^2) = |c|

OpenStudy (zzr0ck3r):

so sqrt(16*c^2) = sqrt(16)sqrt(c^2) = 4sqrt(c^2) = 4|c|

OpenStudy (zzr0ck3r):

does this make sense?

OpenStudy (zzr0ck3r):

@kyleguitarplayer

OpenStudy (anonymous):

Yes it does, keep going please!

OpenStudy (zzr0ck3r):

do this for every term you have 4|c|-|c|+3|c|+3|c| = 9|c|

OpenStudy (zzr0ck3r):

that's it

OpenStudy (anonymous):

4c2 + 5c 4c2 + 7c 8c2 + 5c 8c2 + 7c Which one of those would you think is the right answer?

OpenStudy (zzr0ck3r):

o sorry I read it wrong...

OpenStudy (zzr0ck3r):

sqrt (16c^4) - sqrt (c^2) + 3sqrt (c^2) + sqrt(9c^2) = sqrt(16)*sqrt(c^4) -|c| + 3|c| +3|c| = 4c^2+5|c|

OpenStudy (zzr0ck3r):

they have not told you yet that sqrt(c^2) = |c|, they assume it is sqrt(c^2) = c, this is not really the case yet, but they have not showed you why. so A is the correct answer.

OpenStudy (anonymous):

Thank you so much!